Prove or disaprove .The intersection of two dense subsets of a metric space is also dense in it
1
Expert's answer
2021-07-15T13:11:47-0400

Consider the metric space "\R" of real numbers. The sets "\mathbb Q" of rational numbers and "\R\setminus\mathbb Q" of irrational numbers are dence subsets in "\R," but their intersection "\mathbb Q\cap(\R\setminus\mathbb Q)=\emptyset" is not dence in "\R."

We conclude that in general case the intersection of two dense subsets of a metric space is not dense in it.